# SReferences

### 25 Mar 2024 v1

@@ Abstract
In Figs. S11a-f, we show the spatial distribution of error-mitigated $\langle\hat{Z}_{j}(t)\rangle$ at $t/T=2$, 16, 24, 32, and 40 obtained using ibm_torino for the $L=133$ system with the parameter $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$. These results are compared with those obtained by the MPS method in Figs. S11g-l. Here, we initialise the state $|\psi(0)\rangle$ as a product state with all qubits set to $|0\rangle$, representing the fully polarised state. As shown in Fig. S12c, we also observe a longer-period oscillation in the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ upon introducing non-zero $\theta_{z}$, implying the emergence of a DTQC, despite the initial state differing from that used elsewhere in this study. Interestingly, the period of this oscillation closely resembles that observed in Fig. S7b, where the initial state is set differently. Furthermore, we find in Fig. S12c that the error-mitigated data are in good agreement with the results of MPS over 40-time steps.
@@
Figure S12: (a) Raw data of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}$ for the heavy-hexagonal lattice of $L=133$ qubits with the parameter $(\theta_{x},\theta_{z})=(\pi,0.5\pi)$ obtained using ibm_torino. (b) Same as (a), but with the parameter $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$. (c) Error-mitigated data (yellow diamonds) of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ for the heavy-hexagonal lattice of $L=133$ qubits with the parameter $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$. The results obtained by the MPS method with the bond dimension $\chi=1000$ are also shown with green triangles. Here, the initial state is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
@@
Figures S14b and S14c show the error-unmitigated raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ obtained using ibm_torino for the one-dimensional system of $L=112$ qubits with the parameters $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$ and $(0.8\pi,0.5\pi)$, respectively. Additionally, Fig. S14a shows the error-unmitigated raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ for the same system but with $(\theta_{x},\theta_{z})=(\pi,0.5\pi)$, which are utilised to mitigate errors for $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ with $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$ and $(0.8\pi,0.5\pi)$ shown in Figs. S14b and S14c. From the value of $|\langle\hat{Z}_{\rm avg}(t=80T)\rangle_{0}|$ shown in Fig. S14a, we can estimate the effective quantum circuit volume $v_{\rm eff}\simeq 350$ at $t/T=80$, which is significantly smaller than the quantum circuit volume $v=8880$ (the number of CZ gates in the quantum circuit increases by 111 for each operation of $\hat{U}_{\rm F}$ in the one-dimensional system).
@@
The error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ for the one-dimensional system of $L=112$ qubits with the parameters $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$ and $(0.8\pi,0.5\pi)$ are shown in Fig. S15. In sharp contrast to the case for the heavy-hexagonal lattice of $L=133$ qubits, the period-doubling oscillations observed in the one-dimensional system are much weaker against the deviation of $\theta_{x}$ from $\pi$. We find that the signals with these oscillations decay rather quickly in time steps for other parameter sets of $(\theta_{x},\theta_{z})$ in one dimension, even when $\theta_{z}=0.9$.
Figure S15: Error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ (yellow diamonds) are compared with the results obtained by the MPS simulations with the bond dimensions $\chi=500$ and 1000 (shown as blue crosses and purple diamonds, respectively) for the one-dimensional lattice of $L=112$ qubits. The parameters $(\theta_{x},\theta_{z})$ are (a) $(0.9\pi,0.5\pi)$ and (b) $(0.8\pi,0.5\pi)$.
@@
S5 Tensor network simulations: 2dTNS
@@
[…]
@@
Figure S18: (a) A TNS in the Vidal gauge satisfies the isometric condition after absorbing all but one of their neighboring gauge tensors around the vertex tensors. (b) Deviation from the Vidal gauge is quantified with $\mathcal{C}$. Here, ||$\mathcal{A}\mathclose{\hbox{\set@color${|}$}\kern-1.38889pt\leavevmode\hbox{\set@color${|}$}}|$| represents the Frobenius norm for a tensor $\mathcal{A}$. Each tensor $\mathcal{A}$ after contraction is assumed to be normalised, i.e., $\tilde{\mathcal{A}}=\mathcal{A}/\mathclose{\hbox{\set@color${|}$}\kern-1.38889pt\leavevmode\hbox{\set@color${|}$}}\mathcal{A}\mathclose{\hbox{\set@color${|}$}\kern-1.38889pt\leavevmode\hbox{\set@color${|}$}}$|. (c) Tensor contraction process for evaluating the expectation value $\langle\hat{O}_{i}\rangle$ of a local operator $\hat{O}_{i}$ (denoted as a red square), assuming that a TNS is in the Vidal gauge.

### 12 Mar 2026 v2

@@ Abstract
In Figs. S11a-S11f, we show the spatial distribution of the error-mitigated magnetisation $\langle\hat{Z}_{j}(t)\rangle$ at $t/T=2$, 16, 24, 32, and 40, obtained using ibm_torino for the $L=133$ system with the parameter $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$. Here, the same error mitigation protocol as defined in Eq. (4) of the main text is employed, except that the denominator on the right hand side of Eq. (4) is replaced by the spatial average of the magnetisation over the entire system. These results are compared with those obtained using the MPS method in Figs. S11g-S11l. The initial state $|\psi(0)\rangle$ is prepared as a product state with all qubits set to $|0\rangle$, representing a fully polarised state. As shown in Fig. S12c, we also observe a longer-period oscillation in the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ upon introducing non-zero $\theta_{z}$, indicating an IM-DTC response, despite the initial state differing from that used elsewhere in this study. Interestingly, the period of this oscillation closely resembles that observed in Figs. S7b and S8b, where the initial state is set differently. Furthermore, we find in Fig. S12c that the error-mitigated data are in good agreement with the results of MPS over 40-time steps.
@@
Figure S12: a Raw data of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}$ for the heavy-hexagonal lattice of $L=133$ qubits with the parameter $(\theta_{x},\theta_{z})=(\pi,0.5\pi)$ obtained using ibm_torino. b Same as a, but with the parameter $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$. c Error-mitigated data (yellow diamonds) of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ for the heavy-hexagonal lattice of $L=133$ qubits with the parameter $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$. The results obtained by the MPS method with the bond dimension $\chi=1000$ are also shown with green triangles. Here, the initial state is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
@@
Figures S14b and S14c show the error-unmitigated raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ obtained using ibm_torino for the one-dimensional system of $L=112$ qubits with the parameters $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$ and $(0.8\pi,0.5\pi)$, respectively. Additionally, Fig. S14a shows the error-unmitigated raw data $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ for the same system but with $(\theta_{x},\theta_{z})=(\pi,0.5\pi)$, which are utilised to mitigate errors for $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ with $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$ and $(0.8\pi,0.5\pi)$ shown in Figs. S14b and S14c, respectively. From the value of $|\langle\hat{Z}_{\rm avg}(t=80T)\rangle_{0}|$ shown in Fig. S14a, we can estimate the effective quantum circuit volume $v_{\rm eff}\simeq 350$ at $t/T=80$, which is significantly smaller than the quantum circuit volume $v=8880$ (the number of CZ gates in the quantum circuit increases by 111 for each operation of $\hat{U}_{\rm F}$ in the one-dimensional system).
@@
The error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ for the one-dimensional system of $L=112$ qubits with the parameters $(\theta_{x},\theta_{z})=(0.9\pi,0.5\pi)$ and $(0.8\pi,0.5\pi)$ are shown in Fig. S15. In sharp contrast to the case for the heavy-hexagonal lattice of $L=133$ qubits, the period-doubling oscillations observed in the one-dimensional system are much weaker against the deviation of $\theta_{x}$ from $\pi$. We find that the signals with these oscillations decay rather quickly in time steps for other parameter sets of $(\theta_{x},\theta_{z})$ in one dimension, even when $\theta_{z}=0.9$. These observations are consistent with the expectation that, in one dimension with short-range interactions and without disorder, neither a long-lived DTC nor an IM-DTC response is stabilised.
Figure S15: Error-mitigated data $\langle\hat{Z}_{\rm avg}(t)\rangle$ (yellow diamonds) are compared with the results obtained by the MPS simulations with the bond dimensions $\chi=500$ and 1000 (shown as blue crosses and purple diamonds, respectively) for the one-dimensional lattice of $L=112$ qubits. The parameters $(\theta_{x},\theta_{z})$ are a $(0.9\pi,0.5\pi)$ and b $(0.8\pi,0.5\pi)$.
@@
S5 The role of Ising interaction
In this section, we demonstrate that the Ising interaction $J$ plays an important role in stabilising DTCs. Figure S16 shows the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ with different values of $\theta_{J}$ and $\theta_{x}$, setting $\theta_{z}=0$, for the heavy-hexagonal lattice of $L=28$ qubits (see Fig. 1a), obtained using the state-vector method. Here, we initialise the state $|\psi(0)\rangle$ as a product state with all qubits set to $|0\rangle$ and the magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ is averaged over all qubits. We find that the period-doubling oscillations are quite stable at $\theta_{J}=-0.5\pi$, which is the parameter used in all other cases of this paper and where the DTC and IM-DTC signals are also clearly observed experimentally in ibm_torino. In contrast, for small $\theta_{J}$, magnetisation oscillations are fragile against the deviation of $\theta_{x}$ from $\pi$, leading to a rapid decay of oscillations. This behavior indicates that the DTCs observed on the heavy-hexagonal lattice are supported by strong interactions and are realised within a finite-frequency regime.
Figure S16: Average magnetization $\langle\hat{Z}_{\rm avg}(t)\rangle$ for the heavy-hexagonal lattice of $L=28$ qubits obtained using the state-vector method. The parameter $\theta_{x}$ is set to $0.9\pi$ for ac, $0.85\pi$ for df, and $0.8\pi$ for gi, while the parameter $\theta_{z}$ is fixed at 0. The parameter $\theta_{J}$ of the Ising interaction is $0.125\pi$ in a, d, and g, $0.25\pi$ in b, e, and h, and $0.5\pi$ in c, f, and i. Here, the initial state $|\psi(0)\rangle$ is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state, and the magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ is averaged over all qubits.
S6 Error mitigation based on a global depolarizing noise model
In this section, we provide a detailed description of our error mitigation protocol. When a qubit is depolarised into a fully mixed state $I/2$ with probability $p$, the depolarising noise channel modifies the qubit’s density matrix $\rho$ as [5]
$\rho\rightarrow\mathcal{E}(\rho)=$ $(1-p)\rho+p(I/2).$ (S1)
Accordingly, in an $L$-qubit system, the expectation value of the $i$-th qubit, $\langle\hat{Z}_{i}(t)\rangle_{\text{dp}}$, under depolarizing noise is given by
$\langle\hat{Z}_{i}(t)\rangle_{\text{dp}}\simeq e^{-tp_{i}}\langle\hat{Z}_{i}(t)\rangle,$ (S2)
where $p_{i}$ is the depolarising probability for the $i$-th qubit and $\langle\hat{Z}_{i}(t)\rangle$ represents the magnetisation without quantum noise. The averaged expectation value defined in Eq. (3) of the main text reads
$\langle\hat{Z}_{\rm avg}(t)\rangle_{\text{dp}}=$ $\frac{1}{|A|}\sum_{j\in A}e^{-tp_{j}}\langle\hat{Z}_{j}(t)\rangle$ (S3)
$\simeq$ $\int dpf(p)e^{-tp}\langle\hat{Z}_{\rm avg}(t)\rangle,$ (S4)
where, in the second line, we have introduced a probability density function $f(p)$ to approximate the discrete sum. This approximation assumes a sufficiently large system size and averaging over a large number of qubits. Under these conditions, it is expected that $\langle\hat{Z}_{j}\rangle$ does not depend strongly on each qubit. Assuming a Gaussian distribution $f(p)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left[-\frac{(p-p_{0})^{2}}{2\sigma^{2}}\right]$ with mean $p_{0}$ and variance $\sigma^{2}$, we obtain $\langle\hat{Z}_{\rm avg}(t)\rangle_{\text{dp}}\simeq e^{-tp_{0}}\langle\hat{Z}_{\rm avg}(t)\rangle$. Here, in this study, we adopt a global depolarising noise model, where the depolarisation probability $p$ is nearly uniform across qubits, implying $\sigma\sim 0$. In the following, we argue that this assumption is reasonable under the unitary circuit of the kicked Ising model.
In Fig. S17, we present the depolarisation probability $p_{i}$ for each qubit on ibm_torino, estimated for the kicked Ising model on the heavy-hexagonal lattice of $L=133$ qubits. To estimate $p_{i}$, we fit the function $e^{-tp_{i}}\langle\hat{Z}_{i}(t)\rangle$ to the experimental data $\langle\hat{Z}_{i}(t)\rangle_{\text{dp}}$, based on Eq. (S2). Here, we use the 2dTNS simulation results for the noise-free magnetisation $\langle\hat{Z}_{i}(t)\rangle$ on the right-hand side of Eq. (S2), and the raw experimental data obtained from ibm_torino for $\langle\hat{Z}_{i}(t)\rangle_{\text{dp}}$ on the left-hand side. The experiment was executed on July 29, 2024. Except for several noisy qubits, the values of $p_{i}$ exhibit only weak dependence on the qubit index. This behavior is also reflected in the histograms showing the distribution of $p_{i}$ values. As seen in Figs. S17d–S17f, it is therefore reasonable to approximate the depolarising probabilities by a qubit-independent value $p_{0}$.
Figure S17: Depolarisation probability $p_{i}$ for each qubit on ibm_torino estimated for the kicked Ising model on the heavy-hexagonal lattice of $L=133$ qubits. The model parameter is set to a $\theta_{x}=\pi$, b $0.9\pi$, and c $0.8\pi$, with $\theta_{z}$ fixed at zero. The corresponding values of $p_{0}$ and $\sigma$ are also shown for each $\theta_{x}$. Histograms of $p_{i}$ are shown for d $\theta_{x}=\pi$, e $0.9\pi$, and f $0.8\pi$, where the red dashed lines indicate $p_{0}$.
To further justify our error mitigation protocol based on the global depolarizing noise model, we perform state-vector simulations for the heavy-hexagonal lattice of $L=28$ qubits. In these simulations, quantum noise is incorporated by applying one- and two-qubit depolarising channels to the system’s density matrix $\rho$ as [5, 6]
$\mathcal{E}_{i}^{(1q)}(\rho)=$ $(1-p_{1})\rho+\frac{p_{1}}{3}\sum_{\alpha\neq 0}\sigma_{\alpha,i}\rho\sigma_{\alpha,i}$ (S5)
$\mathcal{E}_{ij}^{(2q)}(\rho)=$ $(1-p_{2})\rho+\frac{p_{2}}{15}\sum_{\alpha,\,\beta}\nolimits^{\prime}\sigma_{\alpha,i}\sigma_{\beta,j}\rho\sigma_{\alpha,i}\sigma_{\beta,j},$ (S6)
where $\sigma_{\alpha,i}$ denotes the Pauli operators acting on qubit $i$, with $\sigma_{\alpha,i}=I$, $X_{i}$, $Y_{i}$, and $Z_{i}$ for $\alpha=0$, 1, 2, and 3, respectively, and the primed sum in Eq. (S6) exclude the identity pair $(\alpha,\beta)=(0,0)$. Each single-qubit gate acting on qubit $i$ is followed by the channel $\mathcal{E}_{i}^{(1q)}$, and each two-qubit gate acting on qubits $i$ and $j$ is followed by $\mathcal{E}_{ij}^{(2q)}$. We assume qubit-independent noise parameters and set $p_{1}=p_{2}/10$, which approximately matches the noise characteristics of the ibm_torino device. Under this noise model, we evaluate the time evolution of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$, as shown in Figs. S18 and  S19 for $p_{2}=0.01$ and $p_{2}=0.005$, respectively. The same error mitigation protocol described in Eq. (4) of the main text is applied but with quantum noises generated according to the depolarising channels defined in Eqs. (S5) and (S6). We average over 100 independent noise realizations. More precisely, for each time step, we take sample averages over 100 noise realizations separately for the numerator and denominator in Eq. (4), using the same set of noise realizations for both. The final error-mitigated value is then obtained by taking the ratio of these sample-averaged quantities. Comparing Figs. S18 and S19 with Fig. S2, we find that $p_{2}=0.01$ is too large to accurately characterise the quantum noises in ibm_torino. On the other hand, the behavior of $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained with $p_{2}=0.005$ (see Fig. S19) closely resembles that observed in the experimental results (see Fig. S2), at least for $t/T<50$. Note that the average error rate of CZ gate on ibm_torino is approximately 0.006.
From the above discussions, we conclude that the global depolarising noise model provides an effective description of quantum noise in the observation of $\langle\hat{Z}_{\rm avg}(t)\rangle$ for the kicked Ising model. The effectiveness of this model for mitigating quantum noise has also been pointed out in Ref. [7]. In quantum simulations of the kicked Ising model with large system size $L$, the number of gates per circuit layer becomes substantial, making the depolarising noise model a suitable effective approximation for characterising the accumulated noise.
Figure S18: Error-mitigated average magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ (yellow diamonds) are compared with the noise-free results (black squares), both obtained by the classical state-vector method, for the heavy-hexagonal lattice of $L=28$ qubits. Quantum noise is introduced via the depolarising channels defined in Eqs. (S5) and (S6), with averages taken over 100 independent noise realizations using $p_{2}=10p_{1}=0.01$. The parameters $(\theta_{x},\theta_{z})$ are a $(0.9\pi,0)$, b $(0.9\pi,0.5\pi)$, c $(0.8\pi,0)$, and d $(0.8\pi,0.5\pi)$. The error bars indicated represent the propagated error in the quantities in the numerator and the denominator of Eq. (4).
Figure S19: Same as Fig. S18, but with $p_{2}=10p_{1}=0.005$.
The depolarising noise model is based on the simplifying assumption that all types of Pauli errors act equivalently. This is one reason why it is often justified in random unitary circuits and/or the deep regions of unitary circuits. In the kicked Ising model, particularly for $\theta_{x}\sim\pi$, the effect of the depolarising noise defined in Eq. (S6) is expected to be anisotropic, assuming that the initial state $|\psi(0)\rangle$ is prepared as a product state in the computational basis. This is because phase-flip errors (Z-type) lead to a smaller decay in $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ than bit-flip errors (X- or Y-type) [6]. Assuming that phase-flip errors have no effect, i.e., $\rho=I_{i}Z_{j}\rho I_{i}Z_{j}=Z_{i}I_{j}\rho Z_{i}I_{j}=Z_{i}Z_{j}\rho Z_{i}Z_{j}$, the two-qubit depolarising channel in Eq. (S6) can be modified as
$\tilde{\mathcal{E}}_{ij}^{(2q)}(\rho)=$ $\left(1-\frac{4}{5}p_{2}\right)\rho+\frac{p_{2}}{15}\sum_{\alpha,\beta}\nolimits^{\prime}\sigma_{\alpha,i}\sigma_{\beta,j}\rho\sigma_{\alpha,i}\sigma_{\beta,j},$ (S7)
which results in a reduced fidelity $\tilde{f}:=1-4p_{2}/5\sim f^{4/5}$, where $f=1-p_{2}$. To correct the underestimated fidelity at $\theta_{x}=\pi$, we can propose the following modification to the error mitigation formula [Eq. (4) in the main text]:
$\langle\hat{Z}_{\rm avg}(t)\rangle\approx\frac{\langle\hat{Z}_{\rm avg}(t)\rangle_{0}}{{|\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}|}^{5/4}}.$ (S8)
In this correction, we neglect single-qubit errors, whose probability $p_{1}$ is smaller than $p_{2}$. Figure S20 shows $\langle\hat{Z}_{\rm avg}(t)\rangle$ obtained using the modified error mitigation protocol based on Eq. (S8). Comparing to the mitigation using Eq. (4) [see Fig. S8], the modified protocol yields slightly better agreement with the 2dTNS results, suggesting that accounting for the anisotropy of Pauli errors can improve mitigation accuracy.
Figure S20: Same as Fig. S8, but using the modified error mitigation protocol defined in Eq. (S8).
S7 Out-of-time-order correlators
In this section, we demonstrate that the DTCs observed in the kicked Ising model on a heavy-hexagonal lattice are realized in a prethermal regime. To this end, we perform state-vector simulations to evaluate the out-of-time-order correlator (OTOC) [8, 9], defined as $\langle\hat{X}_{j}(t)\hat{Z}_{k}\hat{X}_{j}(t)\hat{Z}_{k}\rangle$, for a system of $L=21$ qubits arranged as shown in the inset of Fig. S21d. The results are presented in Figs. S21 and S22. The growth of local operators, i.e., quantum scrambling [10, 11], describes the spreading of quantum information in time. OTOCs are widely used as quantitative probes of such scrambling. At early times, the OTOC takes a value of 1 becuase $\hat{X}_{j}(t)$ and $\hat{Z}_{k}$ are initially nonoverlapping and thus commute. As scrambling progresses and the operators fail to commute, the OTOC decreases and eventually saturates to 0. For $\theta_{x}=0.9\pi$ and $0.8\pi$, the OTOC remains near 1 or exhibits persistent oscillations, as shown in Figs. S21a and S21b, respectively. This behavior indicates the absence of significant quantum scrambling up to $t/T<50$. Since scrambling underpins thermalisation in isolated quantum systems [12, 13], its suppression results in the long-lived stability of the prethermal DTCs observed in this regime. In contrast, for $\theta_{x}=0.7\pi$ and $0.6\pi$, the OTOC decays and saturates to 0, as shown in Figs. S21c and S21d, signaling the onset of scrambling. The decay is faster at $\theta_{x}=0.6\pi$ than at $\theta_{x}=0.7\pi$, where a short-lived prethermal plateau appears around $t/T\sim 20$ (Fig. S21c). To examine the spatial dependence of scrambing, we show colormaps of $\langle\hat{X}_{j}(t)\hat{Z}_{20}\hat{X}_{j}(t)\hat{Z}_{20}\rangle$ at various time steps in Fig. S22. The operator spreading is clearly slower for $\theta_{x}=0.9\pi$ and $0.8\pi$ compared to $\theta_{x}=0.7\pi$. Indeed, for $\theta_{x}=0.9\pi$ and $0.8\pi$, several qubits still exhibit OTOC values close to 1 even at $t/T=50$.
Figure S21: Time evolution of the OTOC $\langle\hat{X}_{j}(t)\hat{Z}_{20}\hat{X}_{j}(t)\hat{Z}_{20}\rangle$ for the kicked Ising model on a heavy-hexagonal lattice of $L=21$ qubits, obtained using the state-vector method. The positions of the qubits $j$ are shown in d. The parameter $\theta_{x}$ is a $0.9\pi$, b $0.8\pi$, c $0.7\pi$, and d $0.6\pi$, with $\theta_{z}$ fixed at zero. Here, the initial state is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
Figure S22: Time evolution of the OTOC $\langle\hat{X}_{j}(t)\hat{Z}_{20}\hat{X}_{j}(t)\hat{Z}_{20}\rangle$ for the kicked Ising model on a heavy-hexagonal lattice of $L=21$ qubits (see the inset of Fig. S21d). The parameter $\theta_{x}$ is af $0.9\pi$, gl $0.8\pi$, and mr $0.7\pi$, with $\theta_{z}$ fixed at zero. Time steps $t/T$ are; 1 for a, g, and m; 10 for b, h, and n; 20 for c, i, and o; 30 for d, j, and p; 40 for e, k, and q; and 50 for f, l, and r. Here, the initial state is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
For comparison, we also present the OTOC $\langle\hat{X}_{j}(t)\hat{Z}_{23}\hat{X}_{j}(t)\hat{Z}_{23}\rangle$ for the kicked Ising model on a square lattice of $L=24$ qubits, arranged as shown in the inset of Fig. S23b. The results, obtained using the state-vector method, are shown in Figs. S23 and S24. We find that the OTOCs on the square lattice saturate to zero even for $\theta_{x}=0.9\pi$, whereas those on the heavy-hexagonal lattice remain far from saturation under the same conditions (see Figs. S21a). This comparison highlights that quantum scrambling proceeds significantly more slowly on the heavy-hexagonal lattice, supporting the emergence of stable prethermal DTCs in this geometry. Notably, for $\theta_{x}=0.9\pi$ and $0.8\pi$, we observe a clear suppression of quantum-information propagation at qubits $j=4$ and $16$. This behaviour arises because symmetry charges accumulate at these coordination-3 qubits, effectively acting as blocking points for information flow (see Ref. [14] for details). In contrast, at $\theta_{x}=0.7\pi$, the entanglement propagation overcomes these blocking sites, allowing information to spread throughout the lattice and leading to full quantum scrambling.
Figure S23: Time evolution of the OTOC $\langle\hat{X}_{j}(t)\hat{Z}_{23}\hat{X}_{j}(t)\hat{Z}_{23}\rangle$ for the kicked Ising model on a square lattice of $L=24$ qubits, obtained using the state-vector method. The positions of the qubits $j$ are shown in b. The parameter $\theta_{x}$ is a $0.9\pi$ and b $0.8\pi$, with $\theta_{z}$ fixed at zero. Here, the initial state is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
Figure S24: Time evolution of the OTOC $\langle\hat{X}_{j}(t)\hat{Z}_{23}\hat{X}_{j}(t)\hat{Z}_{23}\rangle$ for the kicked Ising model on a square lattice of $L=24$ qubits (see Fig. S23b). The parameter $\theta_{x}$ is af $0.9\pi$ and gl $0.8\pi$, with $\theta_{z}$ fixed at zero. Time steps $t/T$ are: 1 for a and g; 10 for b and h; 20 for c and i; 30 for d and j; 40 for e and k; and 50 for f and l. Here, the initial state is prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
S8 Long-time dynamics
Figure S25: Long-time evolution of the magnetisation $|\langle\hat{Z}_{\rm avg}(t)\rangle|$ for the kicked Ising model on a heavy-hexagonal lattice of $L=21$ qubits (see the inset in a), obtained using the state-vector method. The parameter $\theta_{x}$ is a $0.9\pi$, b $0.85\pi$, and c $0.8\pi$, with $\theta_{z}$ fixed at zero. The average is taken over the entire systems, with the initial state prepared as a product state with all qubits set to $|0\rangle$, i.e., a fully-polarised state.
Figure S26: Same as Fig. S25, but for a heavy-hexagonal lattice of $L=28$ qubits (see the inset in a).
In this section, we present the long-time dynamics of magnetisation oscillations in the kicked Ising model on a heavy-hexagonal lattice. It is well known that Floquet systems with a small transverse field perturbation $\theta_{x}\sim\pi$ can exhibit two-step relaxation dynamics [16, 15]. Following an initial rapid relaxation, the system enters a metastable prethermal state characterized by slow energy absorption. Eventually, it relaxes to an infinite-temperature state, as predicted by the Floquet eigenstate thermalization hypothesis (ETH) [17, 18].
Figures S25 and S26 show the absolute value of the magnetisation, $|\langle\hat{Z}_{\text{avg}}(t)\rangle|$, averaged over the entire system, for heavy-hexagonal lattices with $L=21$ qubits (see the inset of Fig. S25a) and $L=28$ qubits (see the inset of Fig. S26a), respectively. These results are obtained using the state-vector method with the initial state $|\psi(0)\rangle$ prepared as a product state with all qubits seto to $|0\rangle$. For the $L=21$ qubit system, prethermal plateau-like structures are observed in the time intervals $10<t/T<200$, $10<t/T<70$, and $10<t/T<30$ for $\theta_{x}=0.9\pi$, $0.85\pi$, and $0.8\pi$, respectively. These plateau structures persist for longer durations in the $L=28$ qubit system, appearing in the intervals $10<t/T<1000$, $10<t/T<200$, and $10<t/T<70$ for the same values of $\theta_{x}$. This system size dependence suggests that prethermal DTCs on the heavy-hexagonal lattice become increasingly robust with larger system sizes.
Together with the behavior of the OTOCs discussed in Sec. S7, these results support the conclusion that prethermal DTCs emerge in the parameter range $0.8\pi\lesssim\theta_{x}<\pi$ on the heavy-hexagonal lattice.
S9 Tensor network simulations: 2dTNS
@@
To apply a two-qubit gate to a gauging TNS with bond dimension $\chi$, we employ the time-evolving block decimation (TEBD) method [19], outlined in Fig. S28. Initially, in step (a), the surrounding gauge tensors and the vertex tensors targeted for the two-qubit gate are contracted. Following this, QR decomposition is performed on the vertex tensor in step (b) to optimise the computational efficiency for the subsequent singular value decomposition (SVD) in step (d). All tensors connected to the two-qubit gate, along with the gauge tensor, are contracted in step (c). SVD is then executed on the resulting tensor in step (d), yielding a total of $2\chi$ singular values. To prevent an increase in bond dimension, only the $\chi$ largest singular values are retained, with the remaining space truncated to dimension $\chi$. Subsequently, this truncated diagonal matrix, composed of the $\chi$ largest singular values, replaces the original gauge tensor, with the neighboring vertex tensors updated, in steps (e)-(g).
@@
However, the TEBD process generally disrupts the Vidal gauge, even when we start with a TNS with the Vidal gauge. A procedure to restore this broken gauge is known as regauging. Particularly, simple and well-known regauging methods include trivial simple update (tSU) [29, 30, 31] and belief propagation [28, 3, 32, 33, 34, 35]. In our study, we adopt the tSU regauging method outlined in Fig. S30. The tSU is a procedure essentially equivalent to that described in Fig. S28, except that the two-qubit gate is replaced with identity. Repeatedly applying this procedure across all edges in the tensor network improves the gauge and ultimately restores the Vidal gauge, in principle. While there is a mathematical proof for this in a tree tensor network [35], it is not established for a general tensor network with a loop structure. In our case, we assess the error $\mathcal{C}$ from the Vidal gauge, as defined in Fig. S29(b), at each time step, and iterate the tSU until it sufficiently converges. Moreover, we observe no significant difference between taking the Vidal gauge at every time step and only before measuring a physical quantity. This suggests that the accuracy of a TNS itself at each time step is not greatly influenced by whether the Vidal gauge is taken or not. Additionally, we performed calculations using the belief propagation [28, 3], and found that the results coincide with those obtained using the tSU, indicating convergence to the same fixed point [35].